OpenStudy (anonymous):

Solve. Express your answer as exact roots (d+1\2)^2+1

OpenStudy (anonymous):

\((d+\frac{1}{2})^2 +1\) First we FOIL: First, Outside, Inside, Last First: \(d*d=d^2\) Outside: \(d * \frac{1}{2} = \frac{d}{2}\) Inside: \(\frac{1}{2} * d = \frac{d}{2}\) Last: \(\frac{1}{2} * \frac{1}{2} = \frac{1}{4}\) Add them all together to get: (hint: don't forget the +1 at the end.) \(d^2 + \frac{d}{2} + \frac{d}{2} + \frac{1}{4} + 1 \) Combine like terms: \(\frac{d}{2} + \frac{d}{2} = \frac{2d}{2} = d\) \(1 + \frac{1}{4} = \frac{4}{4} + \frac{1}{4} = \frac{5}{4}\) \(d^2 + d + \frac{5}{4}\) Now factor or use the quadratic equation: \(\frac{b \pm \sqrt{b^2-4ac}}{2a}\) \(ax^2 +bx + c\) Substitute A, B, and C: \(\frac{1 \pm \sqrt{1^2-4(1)(\frac{5}{4})}}{2(1)}\) (hint: \(4 * 1 * \frac{5}{4} = \frac{20}{4} = 5\) \(\frac{1 \pm \sqrt{1-5}}{2}\) \(\frac{1 \pm \sqrt{-4}}{2}\) \(\frac{1\pm2i}{2}\) \(\frac{1+2i}{2} = \frac{3i}{2}\) or \(\frac{1-2i}{2} = \frac{-1i}{2}\) Not exactly a pre-algebra question, unless you have entered it wrong.